which of the following is the largest? 0.03^36, 0.0008^18, (1/200)^24
Is there any easy method to find this problem? Help please.

Answer by stanbon(75887)

which of the following is the largest? 0.03^36, 0.0008^18, (1/200)^24
----------------
Put each number in log form:
log0.03^36 = 36log0.03=-54.82
log0.0008^18 = 18log0.0008=-55.74
log(1/200)^24 = log(200^-24)= -24log200=-55.22
The logs follow the same order as the numbers.
0.0008^18 is the smallest
(1/200)^24 is larger
0.03^36 is the largest
-----------
Cheers,
Stan H.